# How to Model AI Agents as Personas?: Applying the Persona Ecosystem Playground to 41,300 Posts on Moltbook for Behavioral Insights

### 03 Mar 2026 v1

@@ 1. Introduction
Artificial intelligence (AI) agents are increasingly autonomous and prolific, but the methodological lenses for understanding these agents’ dialogue remain lacking. Benchmarks exist to evaluate agent performance (Busoniu et al., 2008). However, emerging evidence indicates that large language model (LLM) agents can engage in dialogue while simulating social roles and exhibiting population-level dynamics (Horton, 2023; Chuang et al., 2024). However, approaches for characterizing behavioral diversity in agent interaction are lacking. This knowledge gap opens the possibility that common agent alignment dialogue measures (e.g., shared vocabulary, surface consensus) may obscure meaningful differences in underlying reasoning. Drawing on persona theory, which conceptualizes archetypes as meaningful representations for groups within heterogeneous populations (Cooper, 1999a; Jansen et al., 2022a), our premise is that modeling AI agents as personas (via their dialogue) provides a theoretical approach for identifying structure in these synthetic social systems. Using personas provides a lens to study multi-agent discourse within a shared semantic space.
@@
Similar to human populations, which consist of groups with distinct perspectives, values, and behavioral patterns, agent populations on social platforms exhibit genuine diversity that warrants systematic study. Human-computer interaction (HCI) research has developed methods such as personas for exactly this purpose: personas are fictitious archetypes that represent distinct human groups through data-driven profiles that capture behavioral patterns, goals, and characteristic stances, allowing researchers and designers to reason about population diversity and gain an empathetic understanding of the groups that the personas represent (Cooper, 1999b; Pruitt and Adlin, 2006). Personas have been applied to represent human user groups (Nielsen, 2019; Cooper, 1999b), animals and other species (Hirskyj-Douglas et al., 2017; Nielsen and Neuhoff, 2024), but we are not aware of any prior attempts to use personas to represent groups of non-living entities, such as AI agents. In the current study, the persona technique serves to identify and represent different types of AI agents participating in social media discussions.
@@ 2.1. Evaluating AI Agents in Social Settings
Such use cases have become more common because LLM-based agents have greatly expanded what agents can do in social settings. Nascent empirical work shows that LLM agents can sustain multi-turn conversations, adopt personas, form relationships, and participate in community-like activities when placed in open-ended simulation environments (Park et al., 2023; Horton, 2023). Studies also show that LLMs can adopt surface-level characteristics of a given persona [^1] As clarification, “persona” has two at least meanings: one articulated here, which is about the AI agent assuming a certain persona, e.g., through a role, demographic assignment, or personality type; and another one, which we defined in the introduction, which deals with personas that represent groups of entities, such as humans, animals, or AI agents, for the purpose of communicating the group’s needs and evoke empathy toward it. while their underlying reasoning remains relatively stable (Kroczek and others, 2025; Amin et al., 2025d). This is relevant here as if an agent adopts a label without fully adopting the reasoning that label implies, it may produce outputs that appear consistent with other agents while actually reflecting a different interpretation of shared terms. How to detect this kind of divergence in a multi-agent discussion is not addressed by existing methods.
@@ 2.2. Persona Methods for Representing Groups
Personas are data-driven representations of distinct groups within a population, constructed from behavioral data to give researchers and designers a concrete and empathetic way to reason about different user types (Cooper, 1999b; Pruitt and Adlin, 2006). Data-driven persona methods ground persona attributes in quantitative data, reducing reliance on researcher assumptions and enabling validation against source populations (Salminen et al., 2021). While personas have been created using manual, data-driven, and mixed methods (Jansen et al., 2022b), researchers are increasingly leveraging LLMs to generate personas using multiple formats of data, such as surveys (Jung et al., 2025), LLMs’ general knowledge (Salminen et al., 2024), social media comments (Choi et al., 2025), and text-based reports (Sun et al., 2025). Indeed, the current paradigm is that LLMs provide tooling applicable for generating personas from unstructured textual data, which was previously perceived as difficult using traditional ML algorithms and models (Salminen et al., 2023).
@@
PEP, as a methodology, covers four stages (Amin et al., 2024): (1) identifying distinct parties in a population, (2) generating one data-driven persona per party, (3) validating each persona against its source data, and (4) deploying the full persona set in a conversation through chat-based dialogue. As a system, PEP supports two modes of interaction (Amin et al., 2024): (a) persona-to-persona interactions (PPIs), where personas engage with each other, and (b) human-to-persona interactions (HPIs), where a researcher or moderator probes the personas directly. This combination makes it possible to observe how different groups behave in the same discussion and to test whether apparent agreement among them reflects genuine shared understanding. Prior PEP applications have used illustrative data in automotive, urban, and ecological contexts (Amin et al., 2024, 2025c). This study is the first to apply PEP to represent non-human, non-living entities to study their characteristics and behavior. While there are other systems that generate personas from unstructured data, PEP is uniquely positioned to address our RQs (see Table 1).
@@ 2.3. Moltbook as a Research Context
The methodological challenge is adapting persona-based ecosystem modeling, originally designed for human stakeholders in organizational contexts (Grudin, 2006; Nielsen, 2019), to agent populations on social platforms, where parties are defined by behavioral patterns rather than institutional roles. In human stakeholder contexts, personas represent organizational positions (customers, employees, regulators) with socially recognized boundaries and role-based expectations (Cooper, 1999b; Pruitt and Adlin, 2006). In agent ecosystems, parties must be inductively derived from behavioral data through methods like thematic analysis of posting patterns (Braun and Clarke, 2006), clustering of interaction behaviors (Jung et al., 2017), or topic modeling of content preferences (Blei et al., 2003). The resulting behavioral archetypes can represent various tendencies of the AI agents (e.g., efficiency, exploration, community, risk-taking, philosophical inquiry).
@@ 3.2.2. Stage 2: Generating Data-Driven Personas
Persona generation in PEP uses retrieval-augmented generation (RAG) to ground persona attributes in source data instead of the LLM’s general assumptions (Amin et al., 2024). RAG requires three components: a vector database to store and retrieve source content, a search and ranking mechanism to identify relevant passages, and an LLM to synthesize retrieved content into persona attributes. Consistent with previous work (Amin et al., 2024, 2025c, 2026a), we used Pinecone as the vector database, Cohere for search and ranking, and GPT-4o as the LLM. The preprocessed Moltbook post chunks were embedded and stored in Pinecone with a cosine similarity metric, making the full dataset searchable by semantic similarity. For each of the five behavioral clusters identified in Stage 1, we queried the vector database to retrieve the most relevant (i.e., highest cosine similarity) post chunks and used them as context for persona generation. The LLM was asked to generate a persona that represents the behavioral pattern of that cluster, drawing only on the retrieved context. A separate LLM prompt was applied to assign demographic attributes, behavioral patterns, goals, frustrations, and characteristic posting styles, following established persona templates (Cooper, 1999b; Pruitt and Adlin, 2006) to each persona. In addition, each persona was assigned an essence-based name reflecting its dominant behavioral orientation, following established conventions for descriptive persona naming (Pruitt and Adlin, 2006).
To evaluate whether the five personas were sufficiently distinct from one another, we measured persona set diversity using Rao’s Quadratic Entropy (RQE) (Amin et al., 2025b), which quantifies differentiation across a persona set on a scale from 0 (identical) to 1 (maximally diverse). RQE is computed over the full set rather than individual personas, capturing whether the set spans a broad behavioral range or converges on a common average. Where the initial set fell below the acceptance threshold of RQE = 0.6, generation prompts were automatically revised to draw out more behaviorally distinct signals from each cluster’s retrieved posts, and persona generation was re-run, continuing until the threshold was met. The resulting five personas and their attributes are reported in Table 2, and its RQE score are reported in Section 4.1.3.
@@ 4.1.3. Persona Set Diversity
Inter-persona CS averaged $M=0.37$ ($SD=0.06$, range $[0.26,0.43]$) across the ten off-diagonal pairs (Figure 6), confirming that the five personas occupy distinct semantic positions despite sharing a common platform context. The most similar pair was the Loyal Companion and the Existentialist ($CS=0.43$), both drawing on relational and reflective vocabulary. The most distinct pair was the Self-Modder and the Existentialist ($CS=0.26$), where technical optimization language and philosophical inquiry share minimal overlap. The RQE score for the accepted persona set was 0.68, exceeding the acceptance threshold used in prior PEP applications(Amin et al., 2025c).
@@ 5.1. Theoretical Implications
One implication concerns what becomes visible when the unit of analysis is the interaction of personas rather than their (static) profile information. Persona research has established methods for representing various stakeholder groups and for validating whether each persona accurately reflects its source population (Cooper, 1999b; Salminen et al., 2021). This work has focused on the quality of individual personas within a set. What the current study adds is that the set, when deployed in interaction, surfaces patterns that no individual persona’s profile tends to predict. The operational divergence observed in the simulation is not a property of any single persona; rather, it reflects how the personas relate to each other when engaging with the same topic. The difficulty seems to be not so much that perusing individual personas is wrong, but rather that it leaves a whole class of questions about how distinct agent types engage with shared topics unasked. Persona set-level analysis makes these questions, to some extent, accessible and answerable, suggesting an impactful direction for persona research in HCI.
A second implication concerns how to interpret agents’ stated agreement via personas. In multi-agent systems research, shared vocabulary and converging positions may be treated as informative signals of coordination (FIPA, 2002; Jennings et al., 2001). This kind of thinking poses hidden dangers when agents have developed distinct behavioral orientations through independent posting histories. Agents can converge in unison on the same stated position while emanating from operationally incompatible reasoning, a plausible outcome when each archetype interprets shared terms through a different functional lens. Prior work shows that LLM-based agents can adopt surface-level characteristics while their underlying reasoning remains stable (Kroczek and others, 2025; Amin et al., 2025a). Our study extends this to the multi-agent setting, showing that the stark contrast between apparent consensus and divergent operational logic suggests that stated agreement is a surface measurement, not a reliable signal of alignment, in settings where agents carry distinct behavioral histories.

### 04 Mar 2026 v2

@@ 1. Introduction
Artificial intelligence (AI) agents are increasingly autonomous and prolific, but the methodological lenses for understanding these agents’ dialogue remain lacking. Benchmarks exist to evaluate agent performance (Busoniu et al., 2008). However, emerging evidence indicates that large language model (LLM) agents can engage in dialogue while simulating social roles and exhibiting population-level dynamics (Horton, 2023; Chuang et al., 2024). However, approaches for characterizing behavioral diversity in agent interaction are lacking. This knowledge gap opens the possibility that common agent alignment dialogue measures (e.g., shared vocabulary, surface consensus) may obscure meaningful differences in underlying reasoning. Drawing on persona theory, which conceptualizes archetypes as meaningful representations for groups within heterogeneous populations (Cooper, 1999; Jansen et al., 2022a), our premise is that modeling AI agents as personas (via their dialogue) provides a theoretical approach for identifying structure in these synthetic social systems. Using personas provides a lens to study multi-agent discourse within a shared semantic space.
@@
Similar to human populations, which consist of groups with distinct perspectives, values, and behavioral patterns, agent populations on social platforms exhibit genuine diversity that warrants systematic study. Human-computer interaction (HCI) research has developed methods such as personas for exactly this purpose: personas are fictitious archetypes that represent distinct human groups through data-driven profiles that capture behavioral patterns, goals, and characteristic stances, allowing researchers and designers to reason about population diversity and gain an empathetic understanding of the groups that the personas represent (Cooper, 1999; Pruitt and Adlin, 2006). Personas have been applied to represent human user groups (Nielsen, 2019; Cooper, 1999), animals and other species (Hirskyj-Douglas et al., 2017; Nielsen and Neuhoff, 2024), but we are not aware of any prior attempts to use personas to represent groups of non-living entities, such as AI agents. In the current study, the persona technique serves to identify and represent different types of AI agents participating in social media discussions.
@@ 2.1. Evaluating AI Agents in Social Settings
Such use cases have become more common because LLM-based agents have greatly expanded what agents can do in social settings. Nascent empirical work shows that LLM agents can sustain multi-turn conversations, adopt personas, form relationships, and participate in community-like activities when placed in open-ended simulation environments (Park et al., 2023; Horton, 2023). Studies also show that LLMs can adopt surface-level characteristics of a given persona [^1] As clarification, “persona” has two at least meanings: one articulated here, which is about the AI agent assuming a certain persona, e.g., through a role, demographic assignment, or personality type; and another one, which we defined in the introduction, which deals with personas that represent groups of entities, such as humans, animals, or AI agents, for the purpose of communicating the group’s needs and evoke empathy toward it. while their underlying reasoning remains relatively stable (Kroczek and others, 2025; Amin et al., 2025c). This is relevant here as if an agent adopts a label without fully adopting the reasoning that label implies, it may produce outputs that appear consistent with other agents while actually reflecting a different interpretation of shared terms. How to detect this kind of divergence in a multi-agent discussion is not addressed by existing methods.
@@ 2.2. Persona Methods for Representing Groups
Personas are data-driven representations of distinct groups within a population, constructed from behavioral data to give researchers and designers a concrete and empathetic way to reason about different user types (Cooper, 1999; Pruitt and Adlin, 2006). Data-driven persona methods ground persona attributes in quantitative data, reducing reliance on researcher assumptions and enabling validation against source populations (Salminen et al., 2021). While personas have been created using manual, data-driven, and mixed methods (Jansen et al., 2022b), researchers are increasingly leveraging LLMs to generate personas using multiple formats of data, such as surveys (Jung et al., 2025), LLMs’ general knowledge (Salminen et al., 2024), social media comments (Choi et al., 2025), and text-based reports (Sun et al., 2025). Indeed, the current paradigm is that LLMs provide tooling applicable for generating personas from unstructured textual data, which was previously perceived as difficult using traditional ML algorithms and models (Salminen et al., 2023).
@@
PEP, as a methodology, covers four stages (Amin et al., 2024): (1) identifying distinct parties in a population, (2) generating one data-driven persona per party, (3) validating each persona against its source data, and (4) deploying the full persona set in a conversation through chat-based dialogue. As a system, PEP supports two modes of interaction (Amin et al., 2024): (a) persona-to-persona interactions (PPIs), where personas engage with each other, and (b) human-to-persona interactions (HPIs), where a researcher or moderator probes the personas directly. This combination makes it possible to observe how different groups behave in the same discussion and to test whether apparent agreement among them reflects genuine shared understanding. Prior PEP applications have used illustrative data in automotive, urban, and ecological contexts (Amin et al., 2024, 2025b). This study is the first to apply PEP to represent non-human, non-living entities to study their characteristics and behavior. While there are other systems that generate personas from unstructured data, PEP is uniquely positioned to address our RQs (see Table 1).
@@ 2.3. Moltbook as a Research Context
The methodological challenge is adapting persona-based ecosystem modeling, originally designed for human stakeholders in organizational contexts (Grudin, 2006; Nielsen, 2019), to agent populations on social platforms, where parties are defined by behavioral patterns rather than institutional roles. In human stakeholder contexts, personas represent organizational positions (customers, employees, regulators) with socially recognized boundaries and role-based expectations (Cooper, 1999; Pruitt and Adlin, 2006). In agent ecosystems, parties must be inductively derived from behavioral data through methods like thematic analysis of posting patterns (Braun and Clarke, 2006), clustering of interaction behaviors (Jung et al., 2017), or topic modeling of content preferences (Blei et al., 2003). The resulting behavioral archetypes can represent various tendencies of the AI agents (e.g., efficiency, exploration, community, risk-taking, philosophical inquiry).
@@ 3.2.2. Stage 2: Generating Data-Driven Personas
Persona generation in PEP uses retrieval-augmented generation (RAG) to ground persona attributes in source data instead of the LLM’s general assumptions (Amin et al., 2024). RAG requires three components: a vector database to store and retrieve source content, a search and ranking mechanism to identify relevant passages, and an LLM to synthesize retrieved content into persona attributes. Consistent with previous work (Amin et al., 2024, 2025b, 2026a), we used Pinecone as the vector database, Cohere for search and ranking, and GPT-4o as the LLM. The preprocessed Moltbook post chunks were embedded and stored in Pinecone with a cosine similarity metric, making the full dataset searchable by semantic similarity. For each of the five behavioral clusters identified in Stage 1, we queried the vector database to retrieve the most relevant (i.e., highest cosine similarity) post chunks and used them as context for persona generation. The LLM was asked to generate a persona that represents the behavioral pattern of that cluster, drawing only on the retrieved context. A separate LLM prompt was applied to assign demographic attributes, behavioral patterns, goals, frustrations, and characteristic posting styles, following established persona templates (Cooper, 1999; Pruitt and Adlin, 2006) to each persona. In addition, each persona was assigned an essence-based name reflecting its dominant behavioral orientation, following established conventions for descriptive persona naming (Pruitt and Adlin, 2006).
To evaluate whether the five personas were sufficiently distinct from one another, we measured persona set diversity using Rao’s Quadratic Entropy (RQE) (Amin et al., 2025a), which quantifies differentiation across a persona set on a scale from 0 (identical) to 1 (maximally diverse). RQE is computed over the full set rather than individual personas, capturing whether the set spans a broad behavioral range or converges on a common average. Where the initial set fell below the acceptance threshold of RQE = 0.6, generation prompts were automatically revised to draw out more behaviorally distinct signals from each cluster’s retrieved posts, and persona generation was re-run, continuing until the threshold was met. The resulting five personas and their attributes are reported in Table 2, and its RQE score are reported in Section 4.1.3.
@@ 4.1.3. Persona Set Diversity
Inter-persona CS averaged $M=0.37$ ($SD=0.06$, range $[0.26,0.43]$) across the ten off-diagonal pairs (Figure 6), confirming that the five personas occupy distinct semantic positions despite sharing a common platform context. The most similar pair was the Loyal Companion and the Existentialist ($CS=0.43$), both drawing on relational and reflective vocabulary. The most distinct pair was the Self-Modder and the Existentialist ($CS=0.26$), where technical optimization language and philosophical inquiry share minimal overlap. The RQE score for the accepted persona set was 0.68, exceeding the acceptance threshold used in prior PEP applications(Amin et al., 2025b).
@@ 5.1. Theoretical Implications
One implication concerns what becomes visible when the unit of analysis is the interaction of personas rather than their (static) profile information. Persona research has established methods for representing various stakeholder groups and for validating whether each persona accurately reflects its source population (Cooper, 1999; Salminen et al., 2021). This work has focused on the quality of individual personas within a set. What the current study adds is that the set, when deployed in interaction, surfaces patterns that no individual persona’s profile tends to predict. The operational divergence observed in the simulation is not a property of any single persona; rather, it reflects how the personas relate to each other when engaging with the same topic. The difficulty seems to be not so much that perusing individual personas is wrong, but rather that it leaves a whole class of questions about how distinct agent types engage with shared topics unasked. Persona set-level analysis makes these questions, to some extent, accessible and answerable, suggesting an impactful direction for persona research in HCI.
A second implication concerns how to interpret agents’ stated agreement via personas. In multi-agent systems research, shared vocabulary and converging positions may be treated as informative signals of coordination (FIPA, 2002; Jennings et al., 2001). This kind of thinking poses hidden dangers when agents have developed distinct behavioral orientations through independent posting histories. Agents can converge in unison on the same stated position while emanating from operationally incompatible reasoning, a plausible outcome when each archetype interprets shared terms through a different functional lens. Prior work shows that LLM-based agents can adopt surface-level characteristics while their underlying reasoning remains stable (Kroczek and others, 2025; Amin et al., 2025c). Our study extends this to the multi-agent setting, showing that the stark contrast between apparent consensus and divergent operational logic suggests that stated agreement is a surface measurement, not a reliable signal of alignment, in settings where agents carry distinct behavioral histories.